# Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy

### v1

@@ Abstract
Utility and topical relevance are critical measures in information retrieval (IR), reflecting system and user perspectives, respectively. While topical relevance has long been emphasized, utility is a higher standard of relevance and is more useful for facilitating downstream tasks, e.g., in Retrieval-Augmented Generation (RAG). When we incorporate utility judgments into RAG, we realize that the topical relevance, utility, and answering in RAG are closely related to the three types of relevance that Schutz discussed from a philosophical perspective. They are topical relevance, interpretational relevance, and motivational relevance, respectively. Inspired by the dynamic iterations of the three types of relevance, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to promote each step of the cycle of RAG. We conducted extensive experiments on multi-grade passage retrieval and factoid question-answering datasets (i.e., TREC DL, WebAP, and NQ). Experimental results demonstrate significant improvements in utility judgments, ranking of topical relevance, and answer generation upon representative baselines, including multiple single-shot utility judging approaches. Our code and benchmark can be found at https://anonymous.4open.science/r/ITEM-B486/.
Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy
Hengran Zhang1,2 Keping Bi1,2 Jiafeng Guo1,2 Xueqi Cheng1,2 1CAS Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2School of Computer Science and Technology, University of Chinese Academy of Sciences; {zhanghengran22z, bikeping, guojiafeng, cxq}@ict.ac.cn,
@@ 1 Introduction
Utility and topical relevance measure Information Retrieval (IR) performance from the views of users and systems, respectively Bruce (1994). Utility refers to the usefulness of retrieval items to an information seeker, gauged by their “value”, “appropriateness in resolution of a problem” or “reduction of uncertainty”, etc. Saracevic (1996); Schamber and Eisenberg (1988); Saracevic (1975). In contrast, topical relevance measures the matching between the subjects of queries and documents, which indicates aboutness Saracevic et al. (1988). Topical relevance does not necessarily mean utility, while utility indicates a high standard of relevance.
Since topical relevance is relatively easy to observe and measure Schamber et al. (1990), the studies of IR models have been primarily focused on improving the system view of relevance for a long time Bruce (1994). In recent years, most IR models have been experimented on the MSMARCO dataset Nguyen et al. (2016) and optimized towards the binary labels, which is inadequate to capture utility. More recently, leveraging Large Language Models (LLMs) OpenAI (2022); Meta (2024); Jiang et al. (2023a); OpenAI (2023) for topical relevance judgments Khramtsova et al. (2024); Faggioli et al. (2023); Thomas et al. (2024); Sun et al. (2023) has drawn more attention than judging utility Zhang et al. (2024).
@@
In the modern LLM era, Retrieval-Augmented Generation (RAG) has become a hot research topic that facilitates LLMs with external knowledge Xie et al. (2023); Shi et al. (2023); Izacard et al. (2023); Su et al. (2024); Glass et al. (2022). It requests retrieval results useful for answering the question. Given the constrained bandwidth of LLM inputs, it is essential to prioritize high-value results to guide LLMs. Consequently, utility needs to be emphasized more than topical relevance in RAG. In this paper, we aim to promote the utility judgment performance of LLMs so that relevance ranking can be boosted and RAG can be enhanced by high-utility references.
@@
We find that the interpretation of relevance in philosophy by Schutz Schutz (1970) is closely related to RAG. Schutz considered relevance as the property that determines the connections and relations in our lifeworld. He identified three types of basic and interdependent relevance that interact dynamically within a “system of relevancies” Saracevic (1996): (i) Topical relevance, which refers to the perception of what is separated from one’s experience to form one’s present object of concentration; (ii) Interpretational relevance, which involves the past experiences in understanding the currently concerned object; and (iii) Motivational relevance, which refers to the course of action to be adapted based on the interpretations. The motivational relevance, in turn, helps obtain additional materials to become a user’s new experience, which further facilitates topical and interpretational relevance. Schutz posited that one’s perception of the world may be enhanced by this dynamic interaction, as shown in Figure 1. We consider that topical relevance, interpretational relevance, and motivational relevance are closely related to topical relevance, utility, and answering respectively in RAG. Inspired by the dynamic interaction of the relevance system, we believe the dynamic interactions between the three components in RAG can promote the performance of each step.
To verify the idea, we leverage LLMs to perform each step in RAG shown in Figure 1, i.e., the judgments of topical relevance, utility, and answer generation. While previous work Zhang et al. (2024) has shown that utility judgments via LLMs can be enhanced by simultaneously answering the question, iteratively conducting utility judgments and question answering (QA) has not been explored, not to mention that topic relevance can also be involved in the iteration. We propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps.
@@
We experiment on various information-seeking tasks, i.e., multi-grade passage retrieval on TREC DL Craswell et al. (2020), multi-grade non-factoid answer passage retrieval on WebAP Yang et al. (2016), and factoid QA Kwiatkowski et al. (2019). For multi-grade passage retrieval, we consider the ones with the highest grade to be of utility and we focus on the performance of utility judgments and topical relevance ranking. For factoid QA, we emphasize the answer accuracy. Experimental results have demonstrated that ITEM can significantly outperform competitive baselines including various single-shot judgment approaches in terms of utility judgments, ranking of topical relevance, and answer generation, which confirms the viability of adaptation of Schutz’s viewpoint of the relevance system into RAG.
@@ 2 Related Work
Relevance-Focused Information Retrieval. Relevance is a crucial concept in information retrieval (IR), and the automatic generation of relevance judgments has been a long-standing goal in IR that has been studied over the past decades Saracevic (1975); Järvelin and Kekäläinen (2017); Makary et al. (2017, 2016); MacAvaney and Soldaini (2023). Recent studies demonstrate promising outcomes in utilizing LLMs to automatically generate relevance judgments Khramtsova et al. (2024); Faggioli et al. (2023); Thomas et al. (2024) and mainly contain the following three approaches, i.e., pointwise Zhuang et al. (2024); Nogueira and Cho (2019), pairwise Qin et al. (2023); Jiang et al. (2023b), and listwise Sun et al. (2023); Pradeep et al. (2023). For example, Faggioli et al. (2023) used zero/few-shot prompting to instruct LLMs to judge the relevance of an item to a query in a pointwise approach. Sun et al. (2023) analyzed the capabilities of LLMs in the relevance ranking.
@@
Retrieval-Augmented Generation (RAG). RAG approaches are widely employed to mitigate the hallucination issues in large language models (LLMs) Xie et al. (2023); Zhou et al. (2024); Su et al. (2024). The current RAG approaches are categorized as follows: (i) single-round retrieval Borgeaud et al. (2022); Lewis et al. (2020); Glass et al. (2022); Izacard et al. (2023); Shi et al. (2023), which involves using the initial input as a query to retrieve information from an external corpus and then the information is incorporated as part of the input for the model; and (ii) multi-round retrieval Su et al. (2024); Jiang et al. (2023c); Ram et al. (2023); Khandelwal et al. (2020); Trivedi et al. (2023), which need multi-round retrieval based on feedback from LLMs.
@@ 3 Utility Judgments (UJ) via LLMs
We formally define the task of utility judgments via LLMs and propose an Iterative utiliTy judgmEnt fraMework (ITEM) for utility judgments.
@@ 3.1 Notations and Definitions
Given a question and a list of retrieved passages . The goal of utility judgments for LLMs is to identify a set of passages that have utility in answering the question . There are two typical input approaches for LLMs:
Listwise, , where is the instruction of LLMs and is the utility judgments function based on the whole candidate list and the output is a set of passages.
Pointwise, where is the instruction of LLMs and is the utility judgments function with a single passage as the input.
3.2 Sing-Shot Utility Judgments
The most common approach to judge utility is to perform a single-shot utility judgment, i.e., , where is the utility judgments function for LLMs, which can be or . Zhang et al. (2024) proposed to generate a pseudo-answer while conducting utility judgments, which can help LLMs to judge utility better, i.e., .
@@ 3.3 Iterative utiliTy judgmEnt fraMework (ITEM)
Schutz Schutz (1970) emphasized the existence of various types of relevance and underscored the interactivity and interdependence between these various types from a much broader arena than information science. Inspired by the powerful insight, we propose an Iterative utiliTy judgmEnt fraMework (ITEM). Specifically, as shown in Figure 4, dynamic iteration of the topical relevance, pseudo answer generation, and utility judgments.
ITEM with Answering in the Loop (ITEM-A).
In a formal sense, at iteration (), the pseudo answer is generated based on the judging result in the previous iteration and then judged based on the and candidate passages list to get a set of passages that have utility :
@@
where is an answer generation function, is a utility judgment function, which can be or , is the answer prompts for LLMs (as detailed in Figure 2), and is the utility judgments prompts for LLMs (as detailed in Figure 3). can be in two forms: (i) explicit answerto the question ; (ii) implicit answerthat specifies necessary information to answer the question . We consider as the initial candidate set and is the initial results ranked by a retriever like BM25 Robertson et al. (2009).
ITEM with both Answering and Ranking of Topical Relevance in the Loop (ITEM-AR). In Schutz’s theory, topical relevance, motivational relevance, and interpretive relevance are all dynamically affected. In the ITEM-A framework, the topical relevance is not updated during the iteration process. Consequently, we have incorporated a relevance ranking task into the ITEM framework, which ensures that all three tasks are executed in a loop.
@@
Formally, at iteration (), the answer is generated based on the judging result in the previous iteration. Subsequently, given , the passage list is ranked based on the relevance to the question to obtain . Finally, the judging result is derived using the input and the answer :
@@
where and are the relevance ranking function and the relevance ranking prompt for LLMs (as detailed in Figure 5), respectively.
@@
At iteration , we have two ways to produce the set : (i) asking LLMs to identify the set of passages that have utility, which we call ITEM-As or ITEM-ARs; (ii) asking LLMs to provide a ranked passage list with the listwise input approach and we deem the set of first elements in the list as , which we call ITEM-Ar or ITEM-ARr. We set as 5 and more details of selection are shown in Appendix B.4. In our experiments, we find that ITEM-ARr does not improve ranking performance as well as ITEM-Ar (see Section 6.1 for experimental analysis), so we do not use ITEM-ARr in our experiments to obtain .
Full details of all prompts can be found in Appendix A. We stop the iteration when at most iterations are reached or the set of passages considered to have utility does not change, i.e., or . If the value of is relatively small, it may negatively impact the performance of the iteration. Conversely, if the value of is too large, it may increase the computational cost of the iteration. Considering the balance between performance and efficiency, we selected =3. Further discussion about is provided in Section 6.1.
@@ 4.1 Datasets
Our experiments are conducted on three benchmark datasets, including, two retrieval datasets, i.e., TREC DL Craswell et al. (2020), WebAP Yang et al. (2016), and an open-domain question answer (ODQA) dataset, i.e., NQ Kwiatkowski et al. (2019).
TREC. We use the TREC-DL19 and TREC-DL20 datasets Craswell et al. (2020). Judgments of TREC DL are on a four-point scale, i.e., “perfectly relevant”, “highly relevant”, “related”, and “irrelevant”. We consider the passages that are “perfectly relevant” to have utility. We filter questions of two datasets that contain the passages labeled by the highest point and combine them to form a whole dataset, i.e., the TREC dataset in our experiments.
WebAP. WebAP Yang et al. (2016) is a non-factoid answer passage collection built on Gov2. More details about non-factoid answer passage retrieval are provided in Appendix F. It retrieves the top 50 documents with the Sequential Dependency Model (SDM) Metzler and Croft (2005) for the description queries that need to be answered with passages. Relevant documents are annotated, resulting in 3843 passages from 1200 documents, categorized as “perfect”, “excel”, “good”, and “fair”. In our experiments, we split the remaining documents into passages of 2 or 3 sentences to be irrelevant and considered “perfect” to have utility.
NQ. Natural Questions(NQ) consist of factoid questions issued to the Google search engine Kwiatkowski et al. (2019). Each question is annotated with a long answer (typically a paragraph) and a short answer (one or more entities). Following Zhang et al. (2024), we use the questions that have long answers in our experiments. Detailed statistics of all the experimental datasets are shown in Table 16 in the Appendix E.

### v2

# An Iterative Utility Judgment Framework Inspired by Philosophical Relevance via LLMs

@@ Abstract
Relevance and utility are two frequently used measures to evaluate the effectiveness of an information retrieval (IR) system. Relevance emphasizes the aboutness of a result to a query, while utility refers to the result’s usefulness or value to an information seeker. In Retrieval-Augmented Generation (RAG), high-utility results should be prioritized to feed to LLMs due to their limited input bandwidth. Re-examining RAG’s three core components—relevance ranking derived from retrieval models, utility judgments, and answer generation—aligns with Schutz’s philosophical system of relevances, which encompasses three types of relevance representing different levels of human cognition that enhance each other. These three RAG components also reflect three cognitive levels for LLMs in question-answering. Therefore, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to promote each step in RAG. We conducted extensive experiments on retrieval (TREC DL, WebAP), utility judgment task (GTI-NQ), and factoid question-answering (NQ) datasets. Experimental results demonstrate significant improvements of ITEM in utility judgments, ranking, and answer generation upon representative baselines111Our code and benchmark can be found at https://anonymous.4open.science/r/ITEM-B486/..
An Iterative Utility Judgment Framework Inspired by Philosophical Relevance via LLMs
Hengran Zhang1,2,3 Keping Bi1,2,3 Jiafeng Guo1,2, 3 Xueqi Cheng1,2,3 1Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2State Key Laboratory of AI Safety; 3University of Chinese Academy of Sciences; {zhanghengran22z, bikeping, guojiafeng, cxq}@ict.ac.cn,
@@ 1 Introduction
Relevance and utility are two frequently used measures to evaluate Information Retrieval (IR) performance Saracevic (1996, 1975); Saracevic et al. (1988). Relevance usually refers to topical relevance that measures the degree of fit between the subjects of a query and retrieved items, and the criteria of “aboutness” is used Saracevic et al. (1988); Schamber and Eisenberg (1988). In contrast, utility refers to the “usefulness” of retrieval items to an information seeker, of which the criterion is their “value” to the user Saracevic (1996); Saracevic et al. (1988). As the example from the TREC DL dataset shown on Figure 1, topical relevance does not necessarily mean utility, while utility indicates a higher standard of relevance. Since topical relevance is relatively easy to observe and measure Schamber et al. (1990), the studies of IR models have been primarily focused on improving relevance for a long time Bruce (1994).
@@
In the modern LLM era, Retrieval-Augmented Generation (RAG) has become a hot research topic that equips LLMs with external knowledge Xie et al. (2023); Shi et al. (2023); Izacard et al. (2023); Su et al. (2024); Glass et al. (2022). Given the constrained bandwidth of LLM inputs, it is essential to prioritize high-value results to guide LLMs. Consequently, utility needs to be emphasized more than topical relevance in RAG. More recently, Zhang et al. (2024) highlighted the use of LLMs for utility judgments. In this paper, we aim to further promote the utility judgment performance of LLMs so that RAG can be enhanced by high-utility references.
@@
Schutz’s Philosophical Theory of Relevance. Relevance is foundational in information retrieval (IR) and remains widely debated Mizzaro (1998). Saracevic (1996) discussed the nature of relevance in the IR system as the effectiveness of interactive exchange on different levels, and they are non-independent interdependencies, which are primarily influenced by Schutz’s philosophical theory of relevance. Schutz considered relevance as the property that determines the connections and relations in our lifeworld. He identified three types of basic and interdependent relevance that interact dynamically within a “system of relevancies” Saracevic (1996); Schutz (1970): (i) Topical relevance, which refers to the perception of what is separated from one’s experience to form one’s present object of concentration; (ii) Interpretational relevance, which involves the past experiences in understanding the currently concerned object; and (iii) Motivational relevance, which refers to the course of action to be adapted based on the interpretations. The motivational relevance, in turn, helps obtain additional materials to become a user’s new experience, which further facilitates topical and interpretational relevance. Schutz posited that one’s perception of the world may be enhanced by this dynamic interaction, as shown in Figure 2. By incorporating utility judgments into RAG, we can re-examine its three components: topical relevance or relevance ranking derived from retrieval models, utility judgments, and answer generation. Topical relevance is an emerging focus on a topic, utility is the deeper understanding of the topic, and answers indicate the final solution based on the interpretations and will guide users’ actions. Therefore, topic relevance, utility, and derived answer also reflect three cognitive levels for LLMs in question-answering, from low to high, i.e., aboutness, the value for deriving an answer, and the derived answer.
Iterative utiliTy judgmEnt fraMework (ITEM). Inspired by the philosophical theory of relevance, we believe the dynamic interactions between the three components in RAG can promote the performance of each step. To verify the idea, we leverage LLMs to perform each step in RAG shown in Figure 2, i.e., relevance ranking, utility judgments (classification), and answer generation. We propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps. ITEM has two variants depending on whether relevance ranking is involved in the iterations. We are curious to see which option will be better for the tasks: fewer iterations with more components in an iteration, more iterations with fewer components in an iteration, or more iterations with more components.
@@
We experiment on various information-seeking tasks, i.e., multi-grade passage retrieval on TREC DL Craswell et al. (2020), multi-grade non-factoid answer passage retrieval on WebAP Yang et al. (2016), utility judgments benchmark on GTI-NQ Zhang et al. (2024), and factoid QA on NQ Kwiatkowski et al. (2019). For multi-grade passage retrieval, we consider the ones with the highest grade to be of utility and we focus on the performance of utility judgments and topical relevance ranking. For factoid QA, we emphasize the answer accuracy. Experimental results have demonstrated that ITEM can significantly outperform competitive baselines, including various single-shot judgment approaches in terms of utility judgments, ranking of topical relevance, and answer generation, which confirms the viability of adaptation of Schutz’s viewpoint of the relevance system into RAG. We also find that: 1) for difficult tasks (i.e., utility judgments of non-factoid answer passages in WebAP) and complicated candidate passage list (i.e., GTI-NQ), more components in the iteration and multiple iterations are usually more beneficial; 2) for multi-grade relevance ranking tasks, using utility as ranking criterion leads to significantly better multi-grade relevance performance than relevance ranking results even when utility judgment is involved in the iteration; 3) for factoid QA tasks, more iterations with fewer components performs the best, indicating that more components and more iterations are not always needed, especially for simpler tasks.
@@ 2 Related Work
Multi-dimensional relevance. The concept of “relevance” is central to information retrieval theory. Researchers have extensively debated its definition and measurement Mizzaro (1997). Early approaches primarily defined and assessed relevance through exact term matching Vickery (1959) or logical entailment Hillman (1964). However, subsequent empirical studies revealed the limitations of system-oriented relevance analysis, prompting diverse perspectives on relevance Saracevic (1975); Swanson (1986); Saracevic (1996); Lancaster (1968); Goffman and Newill (1964); Kemp (1974); Bruce (1994). For example, Cooper (1971) introduced logical relevance and utility. Saracevic (1996) summarized five frameworks for information science: systems, communication, situational, psychological, and interaction frameworks, and categorized five distinct types of relevance, i.e., 1) system or algorithmic relevance; 2) topical or subject relevance; 3) cognitive relevance or pertinence; 4) situational relevance or utility; and 5) motivational or affective relevance. Bruce (1994) explored cognitive dimensions of relevance. Over time, scholarly consensus has coalesced around two primary perspectives: the system view and user view, with topical relevance and utility serving as their respective representative frameworks.
@@
Retrieval-Augmented Generation (RAG). RAG approaches are widely employed to mitigate the hallucination issues in large language models (LLMs) Xie et al. (2023); Zhou et al. (2024); Su et al. (2024). The current RAG approaches are categorized as follows: (i) single-round retrieval Borgeaud et al. (2022); Lewis et al. (2020); Glass et al. (2022); Izacard et al. (2023); Shi et al. (2023), which involves using the initial input as a query to retrieve information from an external corpus and then the information is incorporated as part of the input for the model; and (ii) multi-round retrieval Su et al. (2024); Jiang et al. (2023b); Ram et al. (2023); Khandelwal et al. (2020); Trivedi et al. (2023), which need multi-round retrieval based on feedback from LLMs.
@@ 3 Utility Judgments (UJ) via LLMs
Schutz’s philosophy of relevances encompasses three types of relevance: topical, interpretational, and motivational relevance, representing different levels of human cognition, and their dynamic interactions can enhance each other. By incorporating utility judgments into RAG and re-examining its three components, i.e., relevance ranking derived from the retrieval models, utility judgments, and answer generation, we realize they closely correspond to Schutz’s philosophical system of relevance. Topic relevance, utility, and derived answer also reflect three cognitive levels for LLMs in question-answering, from low to high, i.e., aboutness, the value for deriving an answer, and the derived answer. Inspired by Schutz’s theory, we presume they can also interact with each other and enhance each other. Therefore, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) for utility judgments.
@@ 3.1 Notations and Definitions
Given a question and a list of retrieved passages based on topical relevance, the goal of utility judgments for LLMs is to identify a set of passages , is the number of passage with utility judged by LLMs. There are two typical input approaches for LLMs to construct the set : pointwise and listwise: The pointwise approach independently evaluates the utility of individual passages, whereas the listwise method assesses the utility of multiple passages with the list input.
3.2 Single-Shot Utility Judgments
The most common approach to judge utility for the is to perform a single-shot utility judgment, i.e., , where is the instruction. Zhang et al. (2024) proposed to generate a pseudo-answer while conducting the utility judgment task, which can help LLMs to judge utility better, i.e., .
@@ 3.3 Iterative utiliTy judgmEnt fraMework (ITEM)
Inspired by Schutz’s theory of relevance in philosophy, we propose an Iterative Utility Judgment Framework (ITEM) for RAG. As illustrated in Figure 3, our framework dynamically iterates among topical relevance ranking, pseudo-answer generation, and utility judgments. We propose two types of loops where two or three components in RAG interact with each other iteratively (ITEM-A and ITEM-AR).
ITEM with Answering in the Loop (ITEM-A). Formally, at each iteration (), given the pseudo answer generated based on the utility judgment result from the previous iteration, we perform utility judgments on the candidate passages list to obtain a set of passages with utility :
@@
where represents the answer prompts for LLMs (as detailed in Figure 4), can be in two forms (details are shown in the Figure 4): (i) explicit answerto the question ; (ii) implicit answerthat specifies the necessary information to answer the question . denotes the utility judgment prompts for LLMs (as detailed in Figure 5). We consider as the initial candidate set, where represents the initial results ranked by a retriever such as BM25 Robertson et al. (2009).
@@
ITEM with both Answering and Ranking of Topical Relevance in the Loop (ITEM-AR). In the ITEM-A framework, topical relevance is not updated during the iteration process. To incorporate dynamic updating of topical relevance, we integrate a relevance ranking task into the ITEM framework, ensuring that all three tasks are executed in a loop. Formally, at iteration (), the answer is generated based on the judging result from the previous iteration. Subsequently, given , the passage list from the previous iteration is ranked based on the relevance to the question, yielding a new ranked list . Finally, the judging result is derived using the ranked list and the answer :
@@
where is the relevance ranking prompt for LLMs (as detailed in Figure 6), respectively.
@@
Overall. At iteration , we have two ways to produce the set : (i) Set-based approach: Asking LLMs to identify the set of passages that have utility using listwise and pointwise input forms, which called ITEM-As or ITEM-ARs variants; (ii) Rank-based approach: Requesting LLMs to provide a ranked passage list based on utility (utility ranking prompt is shown in Appendix H.2) using the listwise input approach and considering the top- passages in the list to build , which called ITEM-Ar or ITEM-ARr variants. We set and more details of are shown in Appendix A.3. We find that ITEM-ARr does not improve ranking performance as well as ITEM-Ar (see Appendix A.4 for experimental analysis), so we do not employ ITEM-ARr in the ranking experiment. The rank-based approach has poor performance on the utility judgment task (details can be found in Appendix A.4), so we only show the performance of the set-based approach on the utility judgment task.
We stop the iteration when at most (=3 in our paper) iterations are reached or the set of selected passages does not change, i.e., or . Full details of all prompts can be found in Appendix H.
@@ 4.1 Datasets
Our experiments are conducted on four benchmark datasets, including two retrieval datasets, i.e., TREC DL Craswell et al. (2020) and WebAP Yang et al. (2016), a utility judgment dataset, i.e., GTI-NQ Zhang et al. (2024), and an open-domain question answer (ODQA) dataset, i.e., NQ Kwiatkowski et al. (2019). Detailed statistics of the experimental datasets are shown in Table Appendix C. We use two representative retrievers to gather candidate passages in for utility judgments on TREC DL, WebAP, and NQ datasets. Construction details can be found in Appendix F.
TREC DL. We use the TREC-DL19 and TREC-DL20 datasets Craswell et al. (2020). Judgments of TREC DL are on a four-point scale, i.e., “perfectly relevant”, “highly relevant”, “related”, and “irrelevant”. We consider the passages that are “perfectly relevant” to have utility. We filter questions of two datasets that contain the passages labeled “perfectly relevant’ and combine them to form a whole dataset, i.e., the TREC DL.
WebAP. WebAP Yang et al. (2016) is a non-factoid answer passage collection built on Gov2. Non-factoid questions usually require longer answers, such as sentence-level or passage-level Keikha et al. (2014a); Yang et al. (2016); Keikha et al. (2014b). Relevant passages are annotated and categorized as “perfect”, “excel”, “good”, and “fair”. Similarly to TREC DL, we considered the “perfect” passages to have utility.
NQ. Natural Questions (NQ) consist of factoid questions issued to the Google search engine Kwiatkowski et al. (2019). Each question is annotated with a long answer (typically a paragraph) and a short answer (one or more entities). Following Zhang et al. (2024), we use the questions that have long answers in our experiments.
GTI-NQ. Ground-truth inclusion (GTI) benchmark is constructed by Zhang et al. (2024) for utility judgment task. The GTI-NQ constructs a candidate passage set of 10 passages for each query sourced from the NQ dataset, comprising the long answer (designated as the utility passage), highly relevant noisy passages, weakly relevant noisy passages, and counterfactual passages.